# Advantage Alignment Algorithms

### v1

@@ Abstract
The growing presence of artificially intelligent agents in everyday decision-making, from LLM assistants to autonomous vehicles, hints at a future in which conflicts may arise from each agent optimizing individual interests. In general-sum games these conflicts are apparent, where naive Reinforcement Learning agents get stuck in Pareto-suboptimal Nash equilibria. Consequently, opponent shaping has been introduced as a method with success at finding socially beneficial equilibria in social dilemmas. In this work, we introduce Advantage Alignment, a family of algorithms derived from first principles that perform opponent shaping efficiently and intuitively. This is achieved by aligning the advantages of conflicting agents in a given game by increasing the probability of mutually-benefiting actions. We prove that existing opponent shaping methods, including LOLA and LOQA, implicitly perform Advantage Alignment. Compared to these works, Advantage Alignment mathematically simplifies the formulation of opponent shaping and seamlessly works for continuous action domains. We also demonstrate the effectiveness of our algorithm in a wide range of social dilemmas, achieving state of the art results in each case, including a social dilemma version of the Negotiation Game.
@@ 1 Introduction
Recent developments in language modeling (GPT) (Radford et al.,, 2018), image synthesis (Diffusion) (Ho et al.,, 2020) and reinforcement learning (Gato) (Reed et al.,, 2022) hint towards a future in which artificially intelligent systems seamlessly integrate in everyday decision making. These systems often act in a decentralized manner and individually optimize for the goals of their users. However, this individual optimization can lead to conflicts, particularly in tasks characterized by having both cooperative and competitive elements. Social dilemmas, introduced by Rapoport and Chammah, (1965) describe this kind of scenario in which each agent acting greedily often leads to an outcome with less utility than that which would have occurred if they had cooperated. A prime example of a social dilemma at a global scale is the climate change problem. Individual and national interests in economic growth often clash with the need for global cooperation to reduce carbon emissions and mitigate environmental degradation. This dichotomy highlights the complexity of aligning individual actions with the collective well-being.
@@
The fast and opaque decision-making processes of machine learning systems make it challenging for humans to supervise every decision. Consequently, there is a pressing need to develop methods that enable agents to autonomously align their interests with each other. Despite this need, the deep reinforcement learning community has traditionally focused its attention on fully cooperative or fully competitive settings, neglecting the nuances of social dilemmas. Sandholm and Crites, (1996) empirically demonstrate that naive reinforcement learning algorithms converge to Pareto suboptimal Nash equilibria in social dilemmas like the Iterated Prisoner’s Dilemma (IPD). To address this gap, opponent shaping algorithms such as Learning with Opponent Learning Awareness (LOLA) (Foerster et al., 2018b, ) have been introduced.
@@
LOLA is an opponent shaping algorithm that directs the behavior of other agents by assuming that they are naive learners and taking gradients with respect to imagined parameter updates. Since then, other opponent shaping algorithms that compute gradients w.r.t. to imaginary parameter updates have demonstrated similar success at partially competitive tasks: SOS (Letcher et al.,, 2021), COLA (Willi et al.,, 2022) and POLA (Zhao et al.,, 2022). More recently LOQA (Aghajohari et al., 2024b, ), proposes and alternative form of opponent shaping by assuming control over the value function of other agents via REINFORCE estimators (Williams,, 1992). This new approach to opponent shaping has significant computational advantages compared to previous methods and establishes the foundation for the Advantage Alignment algorithms.
@@
In this work we introduce Advantage Alignment, a family of algorithms designed with the goal of shaping rational opponents. We do so by making two assumptions about any reinforcement learning algorithm: a reinforcement learning algorithm (1) aims to maximize their own expected return and (2) takes actions proportionally to this expected return. Under these assumptions we demonstrate that opponent shaping reduces to the problem of aligning the advantages of different players and increasing the log probability of an action proportionally to their alignment. We show how this mechanism lies at the heart of some of the existing opponent shaping algorithms, including LOLA and LOQA. By distilling this objective, Advantage Alignment agents are able to shape opponents without imagined parameter updates (LOLA, SOS) or stochastic gradient estimation that relies on automatic differentiation introduced in DiCE (Foerster et al., 2018a, ) (POLA, COLA, LOQA).
Social dilemmas are present in all sorts of human interaction. We are particularly interested in the problem of designing climate negotiation strategies that allow agents to cooperate without centralized authorities. This problem grounds our research to a situation of great concern and applicability for reinforcement learning research, and allows us to identify and overcome key challenges that arise from scale often overlooked in more simple settings like the Iterated Prisoner’s Dilemma (Rapoport and Chammah,, 1965) and the Coin Game (Foerster et al., 2018b, ). With this goal in mind, we apply Advantage Alignment to a continuous variant of the Negotiation Game (also known as the Exchange Game) (Cao et al.,, 2018). In so doing, we aim to begin to address the challenges arising from scale, offering insights and solutions applicable to complex, real-world interactions.
We list our key contributions:
@@
  • We introduce Advantage Alignment and Proximal Advantage Alignment, two opponent shaping algorithms derived from first principles and reliant on policy gradient estimators.
@@
  • We prove that LOLA (and therefore its variations) and LOQA implicitly perform Advantage Alignment through different mechanisms.
@@
  • We extend REINFORCE-based opponent shaping to continuous action environments and achieve state-of-the-art results in a variant of the continuous Negotiation Game (Cao et al.,, 2018).
@@ 2.1 Social Dilemmas
Social dilemmas describe situations in which selfish behavior leads to comparatively poor outcomes for everyone. Such dilemmas are often formalized as normal form games and constitute a subset of general-sum games. A classical example of a social dilemma is the Iterated Prisoner’s Dilemma (IPD) (Rapoport and Chammah,, 1965), in which two players can choose one of two actions: cooperate or defect. In the one-step version of the game, the dilemma occurs because defecting is a dominating strategy, i.e. independently of what the opponent plays the agent is better off playing defect. However, by the reward structure of the game, both the agent and the opponent would achieve a higher utility if they played cooperate simultaneously. Other social dilemmas have been studied in the literature including the Chicken Game and the Coin Game (Lerer and Peysakhovich,, 2018), which has a similar reward structure to IPD but takes place in a grid world. In this paper we introduce a variation of the Negotiation Game (also known as the Exchange Game) (DeVault et al.,, 2015; Lewis et al.,, 2017), with a strong social dilemma component.
@@ 2.3 Reinforcement Learning
$Q^{1}(s_{t},a_{t},b_{t})=r^{1}(s_{t},a_{t},b_{t})+\mathbb{E}_{s_{t+1}}\left[V^{1}(s_{t+1})|s_{t},a_{t},b_{t}\right]$ (2)
@@
$\nabla_{\theta_{1}}V^{1}(\mu)=\mathbb{E}_{\tau\sim\text{Pr}_{\mu}^{\pi^{1},\pi^{2}}}\left[\sum_{t=0}^{\infty}\gamma^{t}A^{1}(s_{t},a_{t},b_{t})\nabla_{\theta_{1}}\log\pi^{1}(a_{t}|s_{t})\right]$ (3)
Here $A^{1}(s,a,b)$ denotes the advantage of the agent taking action $a$ in state $s$ while the opponent takes action $b$.
@@ 3 Opponent Shaping
LOQA (Aghajohari et al., 2024b, ) assumes the opponent’s policy is a softmax over Q-values:
$\hat{\pi}^{2}(b_{t}|s_{t}):=\frac{\exp Q^{2}(s_{t},a_{t},b_{t})}{\sum_{b}\exp Q^{2}(s_{t},a_{t},b)}.$ (4)
The key idea is that these Q-values depend on $\pi^{1}$, and hence the opponent policy $\hat{\pi}^{2}$ can be differentiated w.r.t. $\theta_{1}$:
$\nabla_{\theta_{1}}\hat{\pi}^{2}(b_{t}|s_{t})=\hat{\pi}^{2}(b_{t}|s_{t})\left(\nabla_{\theta_{1}}Q^{2}(s_{t},a_{t},b_{t})-\sum_{b}\hat{\pi}^{2}(b|s_{t})\nabla_{\theta_{1}}Q^{2}(s_{t},a_{t},b)\right).$ (5)
See Appendix A.3 for a derivation of this expression. This dependency of $\pi^{2}$ on $\theta_{1}$ leads to the emergence of an extra term in the policy gradient:
$\nabla_{\theta_{1}}V^{1}(\mu)=\mathbb{E}_{\tau\sim\text{Pr}_{\mu}^{\pi^{1},\pi^{2}}}\left[\sum_{t=0}^{\infty}\gamma^{t}A^{1}(s_{t},a_{t},b_{t})\left(\underbrace{\nabla_{\theta_{1}}\log\pi^{1}(a_{t}|s_{t})}_{\text{policy gradient term}}+\underbrace{\nabla_{\theta_{1}}\log\pi^{2}(b_{t}|s_{t})}_{\text{opponent shaping term}}\right)\right].$ (6)
@@ 4 Advantage Alignment
     Compute agent actor loss, $L_{a}^{1}$, summing equations (3) and (8)
     Compute opponent actor loss, $L_{a}^{2}$, summing equations (3) and (8)
@@
Each opponent $i$ acts proportionally to the exponent of their action-value function: $\pi^{i}(a|s)\propto\exp\left(Q^{i}(s,a)\right)$.
Using equation (6) and substituting $\hat{\pi}^{2}$ in place of $\pi^{2}$ (per assumption 2), we obtain
@@
The first term is the usual policy gradient. The second term is the opponent shaping term and will be our focus. Approximating (5) by ignoring the contribution due to the partition function, this term becomes:
@@
$\mathbb{E}_{\tau\sim\text{Pr}_{\mu}^{\pi^{1},\pi^{2}}}\left[\sum_{t=0}^{\infty}\gamma^{t}A^{1}(s_{t},a_{t},b_{t})\nabla_{\theta^{1}}Q^{2}(s_{t},a_{t},b_{t})\right].$ (7)
@@
The gradient of the Q-value can be estimated by a REINFORCE estimator, which leads to a nested expectation. Aghajohari et al., 2024b showed that this nested expectation can be flattened, allowing efficient estimation from a single trajectory. We take the same approach (see Appendix A.1) to obtain
@@
$\mathbb{E}_{\tau\sim\text{Pr}_{\mu}^{\pi^{1},\pi^{2}}}\left[\sum_{t=0}^{\infty}\sum_{k=t+1}^{\infty}\gamma^{k}A^{1}(s_{t},a_{t},b_{t})A^{2}(s_{k},a_{k},b_{k})\nabla_{\theta^{1}}\text{log }\pi^{1}(a_{k}|s_{k})\right].$ (8)
@@
This is the main result of this work. The expression above captures the essence of opponent shaping: an agent should align its advantages with those of its opponent in order to steer towards trajectories that are mutually beneficial. More precisely, an agent increases the probability of actions that have high product between the sum of its past advantages and the advantages of the opponent at the current time step. This expression for the Advantage Alignment formula is detailed in Appendix A.5. Equation (8) depends only on the log probabilities of the agent, which allows us to create a proximal surrogate objective that closely follows the PPO (Schulman et al., 2017b, ) formulation:
$\mathbb{E}_{\tau\sim\text{Pr}_{\mu}^{\pi^{1},\pi^{2}}}\left[\sum_{t=0}^{\infty}\sum_{k=t+1}^{\infty}\gamma^{k}\min\left\{r_{n}(\theta_{1})A^{1}_{t}A^{2}_{k},\;\text{clip}\left(r_{n}(\theta_{1});1-\epsilon,1+\epsilon\right)A^{1}_{t}A^{2}_{k}\right\}\right]$ (9)
@@
Where $r_{n}$ denotes the ratio between the policy (current) after $n$ updates and the original policy (old). We also denote $A^{i}_{t}:=A^{i}(s_{t},a_{t},b_{t})$, the advantage of agent $i$ at timestep $t$. This surrogate objective in equation (9) is used to formulate the Proximal Advantage Alignment (Proximal AdAlign) algorithm (see appendix A.6 for implementation details).
@@
Why is Assumption 1 necessary? Assumption 1 allows the agent to influence the learning dynamics of the opponent, by controlling the values for different actions. After one iteration of the algorithm the agent changes the Q-values of the opponent for different actions and, since the opponent aims to maximize their expected return (under the new agent policy), it must change its behavior accordingly.
Table 1: Update direction as a function of signs of advantages
$A^{1}$ $A^{2}$ $\boldsymbol{+}$${-}$
$+\boldsymbol{+}$$\boldsymbol{+}$${-}$
$\boldsymbol{-}$${-}$ $\boldsymbol{+}$
4.1 Interpreting Advantage Alignment
Equation (8) yields four possible different cases for controlling the direction of the gradient of the log probability of the policy. As with the usual policy gradient estimator, the sign multiplying the log probability indicates whether the probability of taking an action should increase or decrease. In Table 1 we show the four different quadrants corresponding to the different signs the two advantages can take. We label and interpret the meaning of each of the four quadrants.
Cooperative: The sign is positive since the two advantages (for the agent and opponent) are positive. Therefore it is rational to increase the probability of this action, since by doing so the agent increases the probability of observing trajectories that are mutually beneficial for it and the opponent.
Empathetic: The sign is negative since the action diminishes the opponent’s return. The agent reduces the probability of taking this action because it hurts the opponent, despite the action being beneficial for the agent, hence the name empathetic.
Vengeful: The sign is negative since the action diminishes the agent’s return. The agent reduces the probability of taking this action because it hurts it, despite the action being beneficial for the opponent, hence the name vengeful.
Spiteful: When both advantages are negative, advantage alignment will increase the probability of taking this action despite hurting both agents. This behavior is counterintuitive; we believe it serves as a deterrent, much like mutually assured destruction.
We now relate existing opponent shaping algorithms to advantage alignment, and argue that these algorithms embody these four behaviors. Theorem 1 shows how opponent shaping algorithms derived from LOLA implicitly do advantage multiplication, which lies at the heart of advantage alignment. Theorem 2 proves that LOQA’s opponent shaping term has the same form as that of Advantage Alignment, differing only by a scalar term. Table 1 then holds for algorithms like LOLA, POLA, COLA, SOS and LOQA.

### v3

@@ Abstract
Artificially intelligent agents are increasingly being integrated into human decision-making: from large language model (LLM) assistants to autonomous vehicles. These systems often optimize their individual objective, leading to conflicts, particularly in general-sum games where naive reinforcement learning agents empirically converge to Pareto-suboptimal Nash equilibria. To address this issue, opponent shaping has emerged as a paradigm for finding socially beneficial equilibria in general-sum games. In this work, we introduce Advantage Alignment, a family of algorithms derived from first principles that perform opponent shaping efficiently and intuitively. We achieve this by aligning the advantages of interacting agents, increasing the probability of mutually beneficial actions when their interaction has been positive. We prove that existing opponent shaping methods implicitly perform Advantage Alignment. Compared to these methods, Advantage Alignment simplifies the mathematical formulation of opponent shaping, reduces the computational burden and extends to continuous action domains. We demonstrate the effectiveness of our algorithms across a range of social dilemmas, achieving state-of-the-art cooperation and robustness against exploitation.
@@ 1 Introduction
Recent advancements in artificial intelligence, such as language models like GPT (Radford et al.,, 2018), image synthesis with diffusion models (Ho et al.,, 2020), and generalist agents like Gato (Reed et al.,, 2022), suggest a future where AI systems seamlessly integrate into everyday human decision-making. While these systems often optimize for the goals of their individual users, this can lead to conflicts, especially in tasks that involve both cooperative and competitive elements. Social dilemmas, as introduced by Rapoport and Chammah, (1965), describe scenarios where agents acting selfishly achieve worse outcomes than if they had cooperated. A global example is the climate change problem, where individual and national interests in economic growth often clash with the need for collective action to reduce carbon emissions and mitigate environmental degradation. The challenges we have faced in tackling this problem highlight the complexity of aligning individual interests with collective well-being.
@@
As artificially intelligent systems become ubiquitous, there is a pressing need to develop methods that enable agents to autonomously align their interests with one another. Despite this, the deep reinforcement learning community has traditionally focused on fully cooperative or fully competitive settings, often neglecting the nuances of social dilemmas. Sandholm and Crites, (1996) empirically demonstrated that naive reinforcement learning algorithms tend to converge to the worst Pareto suboptimal Nash equilibria of Always Defect in social dilemmas like the Iterated Prisoner’s Dilemma (IPD). Foerster et al., 2018b , demonstrated that the same is true for policy gradient methods, and introduced opponent shaping (LOLA) to address this gap.
@@
LOLA is an opponent shaping algorithm that influences the behavior of other agents by assuming they are naive learners and taking gradients with respect to simulated parameter updates. Following this approach, other opponent shaping algorithms that compute gradients with respect to simulated parameter updates have shown success in partially competitive tasks, including SOS (Letcher et al.,, 2021), COLA (Willi et al.,, 2022), and POLA (Zhao et al.,, 2022). More recently, LOQA (Aghajohari et al., 2024b, ) proposed an alternative form of opponent shaping by assuming control over the value function of other agents via REINFORCE estimators (Williams,, 1992). This new approach to opponent shaping offers significant computational advantages over previous methods and lays the foundation for our work.
@@
We introduce Advantage Alignment, a family of algorithms designed to shape rational opponents by aligning their advantages when their historic interactions have been positive. We make two key assumptions about reinforcement learning agents: (1) they aim to maximize their own expected return, and (2) they take actions proportionally to this expected return. Under these assumptions, we demonstrate that opponent shaping reduces to aligning the advantages of different players and increasing the log probability of an action proportionally to their alignment. We show that this mechanism lies at the heart of existing opponent shaping algorithms, including LOLA and LOQA. By distilling this objective, Advantage Alignment agents can shape opponents without relying on imagined parameter updates (as in LOLA and SOS) or stochastic gradient estimation that relies on automatic differentiation introduced in DiCE (Foerster et al., 2018a, ) (as in POLA, COLA, and LOQA). Furthermore, we demonstrate that Advantage Alignment preserves Nash Equilibria, ensuring that our algorithms maintain stable strategic outcomes.
We also introduce Proximal Advantage Alignment, which formulates Advantage Alignment as a modification to the advantage function used in policy gradient updates. By integrating this modified advantage into the Proximal Policy Optimization (PPO) (Schulman et al., 2017b, ) surrogate objective, we develop a scalable and efficient opponent shaping algorithm suitable for more complex environments. To identify and overcome challenges that arise from scale—which are often overlooked in simpler settings like the Iterated Prisoner’s Dilemma (Rapoport and Chammah,, 1965) and the Coin Game (Foerster et al., 2018b, )—we apply Advantage Alignment to a continuous variant of the Negotiation Game (Cao et al.,, 2018) and Melting Pot’s Commons Harvest Open (Agapiou et al.,, 2023). In doing so, we aim to demonstrate the scalability of our methods and offer insights and solutions applicable to complex, real-world agent interactions.
Our key contributions are:
@@
  • We introduce Advantage Alignment and Proximal Advantage Alignment (PAA), two opponent shaping algorithms derived from first principles and based on policy gradient estimators.
@@
  • We prove that LOLA (and its variations) and LOQA implicitly perform Advantage Alignment through different mechanisms.
@@
  • We extend REINFORCE-based opponent shaping to continuous action environments and achieve state-of-the-art results in a continuous action variant of the Negotiation Game (Cao et al.,, 2018).
  • We apply PAA to the Commons Harvest Open environment in Melting Pot 2.0 (Agapiou et al.,, 2023), a high dimensional version of the tragedy of the commons social dilemma, achieving state-of-the-art results and showcasing the scalability and effectiveness of our methods.
@@ 2.1 Social Dilemmas
Social dilemmas describe situations in which selfish behavior leads to sub-optimal collective outcomes. Such dilemmas are often formalized as normal form games and constitute a subset of general-sum games. A classical example of a social dilemma is the Iterated Prisoner’s Dilemma (IPD) (Rapoport and Chammah,, 1965), in which two players can choose one of two actions: cooperate or defect. In the one-step version of the game, the dilemma occurs because defecting is a dominant strategy, i.e., independently of what the opponent plays the agent is better off playing defect. However, by the reward structure of the game, both the agent and the opponent would achieve a higher utility if they played cooperate simultaneously. Beyond the IPD, other social dilemmas have been extensively studied in the literature, including the Chicken Game and the Coin Game (Lerer and Peysakhovich,, 2018), the latter of which has a similar reward structure to IPD but takes place in a grid world. In this paper we introduce a variation of the Negotiation Game (also known as the Exchange Game) (DeVault et al.,, 2015; Lewis et al.,, 2017), with a strong social dilemma component. Additionally, we evaluate our method on the Commons Harvest Open environment in Melting Pot 2.0 (Agapiou et al.,, 2023), which exemplifies a large-scale social dilemma. In this environment, agents must balance short-term personal gains from overharvesting common resources against the long-term collective benefit of sustainable use.
@@ 2.3 Reinforcement Learning
$Q^{1}(s_{t},a_{t},b_{t})=r^{1}(s_{t},a_{t},b_{t})+\gamma\cdot\mathbb{E}_{s_{t+1}}\left[V^{1}(s_{t+1})|s_{t},a_{t},b_{t}\right].$ (2)
@@
$\nabla_{\theta_{1}}V^{1}(\mu)=\mathbb{E}_{\tau\sim\text{Pr}_{\mu}^{\pi^{1},\pi^{2}}}\left[\sum_{t=0}^{\infty}\gamma^{t}A^{1}(s_{t},a_{t},b_{t})\nabla_{\theta_{1}}\log\pi^{1}(a_{t}|s_{t})\right].$ (3)
Here $A^{1}(s,a,b):=Q^{1}(s,a,b)-V^{1}(s)$ denotes the advantage of the agent taking action $a$ in state $s$ while the opponent takes action $b$.
@@ 3 Opponent Shaping
LOQA (Aghajohari et al., 2024b, ) performs opponent shaping by controlling the Q-values of the opponent for different actions assuming that the opponent’s policy is a softmax over these Q-values:
$\hat{\pi}^{2}(b_{t}|s_{t}):=\frac{\exp Q^{2}(s_{t},b_{t})}{\sum_{b}\exp Q^{2}(s_{t},b)},$ (4)
where $Q^{2}(s_{t},b_{t}):=\mathbb{E}_{a\sim\pi^{1}}[Q^{2}(s_{t},a,b_{t})]$. The key idea is that these Q-values depend on $\pi^{1}$, and hence the opponent policy $\hat{\pi}^{2}$ can be differentiated w.r.t. $\theta_{1}$: [^1] See Appendix A.3 for a derivation of this expression.
$\nabla_{\theta_{1}}\hat{\pi}^{2}(b_{t}|s_{t})=\hat{\pi}^{2}(b_{t}|s_{t})\left(\nabla_{\theta_{1}}Q^{2}(s_{t},b_{t})-\sum_{b}\hat{\pi}^{2}(b|s_{t})\nabla_{\theta_{1}}Q^{2}(s_{t},b)\right).$ (5)
This dependency of $\pi^{2}$ on $\theta_{1}$ leads to the emergence of an extra term in the policy gradient:
$\nabla_{\theta_{1}}V^{1}(\mu)=\mathbb{E}_{\tau\sim\text{Pr}_{\mu}^{\pi^{1},\pi^{2}}}\left[\sum_{t=0}^{\infty}\gamma^{t}A^{1}(s_{t},a_{t},b_{t})\left(\underbrace{\nabla_{\theta_{1}}\log\pi^{1}(a_{t}|s_{t})}_{\text{policy gradient term}}+\underbrace{\nabla_{\theta_{1}}\log\hat{\pi}^{2}(b_{t}|s_{t})}_{\text{opponent shaping term}}\right)\right].$ (6)
@@ 4.1 Method Description
     Compute agent actor loss, $L_{a}^{1}$, summing equation 3 and equation 8
     Compute opponent actor loss, $L_{a}^{2}$, summing equation 3 and equation 8
@@
Each opponent $i$ acts proportionally to the exponent of their action-value function: $\pi^{i}(a|s)\propto\exp\left(\beta\cdot Q^{i}(s,a)\right)$.
Using Equation 6 and substituting $\hat{\pi}^{2}$ in place of $\pi^{2}$ (per Assumption 2), we obtain:
@@
The first term is the usual policy gradient. The second term is the opponent shaping term and will be our focus. Approximating the opponent’s policy (right hand side of equation 4) by ignoring the contribution due to the partition function, the opponent shaping term becomes:
@@
$\beta\cdot\mathbb{E}_{\tau\sim\text{Pr}_{\mu}^{\pi^{1},\pi^{2}}}\left[\sum_{t=0}^{\infty}\gamma^{t}A^{1}(s_{t},a_{t},b_{t})\nabla_{\theta^{1}}Q^{2}(s_{t},b_{t})\right].$ (7)
@@
The gradient of the Q-value can be estimated by a REINFORCE estimator, which leads to a nested expectation. Aghajohari et al., 2024b empirically showed that this nested expectation can be efficiently estimated from a single trajectory. We take the same approach (see Appendix A.1) to obtain:
@@
$\beta\cdot\mathbb{E}_{\tau\sim\text{Pr}_{\mu}^{\pi^{1},\pi^{2}}}\left[\sum_{t=0}^{\infty}\gamma^{t+1}\left(\sum_{k<t}\gamma^{t-k}A^{1}(s_{k},a_{k},b_{k})\right)A^{2}(s_{t},a_{t},b_{t})\nabla_{\theta^{1}}\text{log }\pi^{1}(a_{t}|s_{t})\right].$ (8)
@@
The expression above captures the essence of opponent shaping: an agent should align its advantages with those of its opponent in order to steer towards trajectories that are mutually beneficial. More precisely, an agent increases the probability of actions that have high product between the sum of its past advantages and the advantages of the opponent at the current time step. For implementation details of the Advantage Alignment formula see Appendix A.6. Equation 8 depends only on the log probabilities of the agent, which allows us to create a proximal surrogate objective that closely follows the PPO (Schulman et al., 2017b, ) formulation:
$\mathbb{E}_{\tau\sim\text{Pr}_{\mu}^{\pi^{1},\pi^{2}}}\left[\min\left\{r_{n}(\theta_{1})A^{*}(s_{t},a_{t},b_{t}),\;\text{clip}\left(r_{n}(\theta_{1});1-\epsilon,1+\epsilon\right)A^{*}(s_{t},a_{t},b_{t})\right\}\right],\$ (9)
where $r_{n}$ denotes the ratio of the policy (new) after $n$ updates of the original policy (old), and:
$A^{*}(s_{t},a_{t},b_{t})=\left(A^{1}(s_{t},a_{t},b_{t})+\beta\gamma\cdot\left(\sum_{k<t}\gamma^{t-k}A^{1}(s_{k},a_{k},b_{k})\right)A^{2}(s_{t},a_{t},b_{t})\right).$ (10)
@@
This surrogate objective in equation 9 is used to formulate the Proximal Advantage Alignment (Proximal AdAlign) algorithm (see appendix A.7 for implementation details).
@@
Why is Assumption 1 necessary? Assumption 1 allows the agent to influence the learning dynamics of the opponent, by controlling the values for different actions. After one iteration of the algorithm the agent changes the Q-values of the opponent for different actions and, since the opponent aims to maximize their expected return, it must change its behavior accordingly.
4.2 Analyzing Advantage Alignment
[…]
Figure 1: (a) The sign of the product of the gamma-discounted past advantages for the agent, and the current advantage of the opponent, indicates whether the probability of taking an action should increase or decrease. (b) The empirical probability of cooperation of Advantage Alignment for each previous combination of actions in the one step history Iterated Prisoner’s Dilemma, closely resembles tit-for-tat. Results are averaged over 10 random seeds, the black whiskers show one std.
Equation 8 yields four possible different cases for controlling the direction of the gradient of the log probability of the policy. As with the usual policy gradient estimator, the sign multiplying the log probability indicates whether the probability of taking an action should increase or decrease. Intuitively, when the interaction with the opponent has been positive (blue in figure 1(a)) the advantages of the agent align with that of the opponent: the advantage alignment term increases the log probability of taking an action if the advantage of the opponent is positive and decreases it if it is negative. In contrast, if the interaction has been negative (red in figure 1(a)) the advantages are at odds with each other: the advantage alignment term decreases the log probability of taking an action if the advantage of the opponent is positive and increases it if it is negative. We now relate existing opponent shaping algorithms to advantage alignment, and argue that these algorithms use the same underlying mechanisms. Theorem 1 shows that LOLA update from Foerster et al., 2018b can be written as a policy gradient method with an opponent shaping term similar to equation 10. This shows the fundamental relationship between opponent-shaping dynamics and advantage multiplications. Theorem 2 proves that LOQA’s opponent shaping term has the same form as that of Advantage Alignment, differing only by a scalar term.